I have been working on solving linear systems I cannot figure out where to move and mutilply the numbers my teacher taught us two ways combination and subsitution.here is one promblem and i was wondering if you could explain where to move the numbers and things because our teacher is of no help in this area and doesnt seem like she ever will be. 4x+y=-1
-5x-y= 0

Answer by askme(9)

since you have two variables (x and y), you need two equations so that you can substitute in order to find each.


step 1: set to linear format y = mx + b
1st equation: y = -4x - 1
2nd equation: -y = 5x + 0 and we will need to get rid of the negative sign so multiply the whole equation by a negative. we do not need to write the zero so we will leave it out.
- (-y = 5x) --> y = -5x


step 2: substitute your y from your 2nd equation (y=-5x) into the y of your 1st equation (y=-4x-1):
-5x = -4x - 1
*notice how there are no more y's in your equation and we combined your two equations into one equation.


step 3: bring the x's to one side and your constants to the other.
-5x = -4x - 1
-5x + 4x = -1
-1x = -1
-1x/-1 = -1/-1
x = 1


step 4: subsitute your x into your original equations. we will use the 1st one:
4x + y = -1
*remember we found that x = 1
4(1) + y = -1
y = -1 - 4
y = -5
step 5: check your work by substituting your x value and y value into the original equation.
4x + y = -1
*remember x=1 and y=-5
4(1) + (-5) = -1
4 - 5 = -1 and your answer checks out therefore the values of your x and y are correct.